Edwin, thank you for your time. Yes, I do need you to work this problem out all the way. X minus 33 1/3% equals 30.
Answer by AnlytcPhil(1116)

Edwin, thank you for your time. Yes, I do need you to work 
this problem out all the way. X minus 33 1/3% equals 30. 

First change the percent 33 1/3% to a fraction:

33 1/3% = (33 1/3)÷100 = (100/3)÷100 = 

       (100/3)÷(100/1) = (100/3)×(1/100) = 1/3

So

X - (33 1/3% of X) = 30

        X - (1/3)X = 30

Multiply every term by 3

  (3)X - (3)(1/3)X = (3)30

Simplify
           3X - 1X = 90 

Combine like terms

                2X = 90

Divide both sides by 2

              2X/2 = 90/2

                 X = 45

Edwin